OpenStudy (anonymous):

Solve. 8 = 2^(x + 4)

OpenStudy (anonymous):

8 = 2^x + 4 2^x = 4 2^x = 2^2 x = 2

OpenStudy (anonymous):

Here are the answer choices: x = -1 x = -4 x = 0 x = 7

OpenStudy (anonymous):

it =0 my friend

OpenStudy (anonymous):

2(x+4)=8 x+4 = 8/2 = 4 x=4-4 = 0

OpenStudy (anonymous):

Great, thank you! :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

itz -1 8 =\[2^{x+4}\] =>\[2^{3}\]=\[2^{x+4}\] therefore 3=x+4 =>x=-1

OpenStudy (jack1):

@alex8ac for 8 = 2^(x + 4) not havin a go at u man, just helpin u out with the symbols we use here the ^ usually means "to the power of" so 8 = 2^(x + 4) in this case means \[8 = 2^{(x + 4) }\] so as @Rex831 said before, you'd either use known powers or logarithms to work this one out \[\log_{2} 8 = \log_{2} 2^{(x+4)}\] so 3 = x +4 so x = -1